# Data Efficient Prediction of excited-state properties using Quantum Neural Networks

### v1

@@ III.1 Datasets
III.2 Model Details
For the siQNN-NN, we use an NN with a single hidden layer comprising 41 parameters with a “tanh” activation function for the hidden layer. The NN is implemented using PyTorch [28], while the QNN is implemented in sQUlearn [29] and Qiskit, and then integrated into the model and trained using PennyLane [30] with state vector simulations. Pre-training is performed with a large learning rate of 0.5 and a maximum of 1,000 iterations to broadly explore the optimization landscape and prevent getting stuck in local minima. The subsequent training of the siQNN-NN is conducted with a learning rate between 0.005 and 0.05 and a maximum of 2,000 iterations to enable smooth convergence. The performance of the model learning on the ground state derived after pre-training labeled siQNN is also compared to the performance of the classical models in the following.
We compare the performance of these models to a purely classical NN that has the same structure as the siQNN-NN, but without the quantum component. The model is trained on the reduced feature space $\overline{\mathcal{X}}$. To ensure comparability, the number of parameters in the NN is equal to the total number of parameters in the siQNN-NN (the sum of parameters in the QNN and the NN). This model is trained until convergence or until the target training loss $\mathcal{L}_{\rm t}$ is reached. Additionally, a support vector regressor (SVR) and a Gaussian process regressor (GPR) with a radial basis function kernel are implemented using scikit-learn [31]. Their hyperparameters are optimized using a grid search, where the performance of the hyperparameters is evaluated on the training dataset.
[…]
@@ III.3 Benchmark
To draw conclusions about the perfomance of the models based on training dataset size independent of the specific choice of points, we sample multiple training datasets $\mathcal{D}_{l}^{L}\subset\mathcal{D}$. Each set $\mathcal{D}_{l}^{L}$ contains $L$ data points. For each $L$, we sample 50 independent datasets, i.e., $l=1,\dots,50$. As a test set, we choose the remaining data points for each $l$ and $L$, i.e., $\mathcal{D}\setminus\mathcal{D}_{l}^{L}$. In the following, all MSE scores shown are derived from the test dataset. We aim to simulate a realistic scenario where sampling is expensive, and the approximate locations of three different regions (attractive, equilibrium, and repulsive) along the dissociation curve are known. From each of these regions one data point is randomly chosen to account for the different behaviors in these regions. These three data points are then used as initial points for a Bayesian optimization, which determines the final data points in the training dataset $\mathcal{D}_{l}^{L}$. However, using a less informed sampling method, e.g., equally spaced training data points from random subsets of the given dataset yields similar results.
@@ III.3.1 LiH
Figure 6 (a) exemplarily shows a boxplot of the MSE for various models for the first transition energy $\Delta E_{T_{1}}$. The median MSE on the test set for the quantum models (siQNN and siQNN-NN) are similar to each other and nearly an order of magnitude lower than those of the classical models (SVR, GPR, NN), which themselves perform similarly across all training dataset sizes considered. Additionally, the variation of the MSE for the siQNN and siQNN-NN is significantly larger than that of the classical models. This is because, for some training dataset samples, the training points poorly represent the overall shape of the target function, causing all models to perform equally poorly. For other training data point combinations, the siQNN and siQNN-NN outperform the classical models by about an order of magnitude in MSE. In cases where other transition energies than $\Delta E_{T_{1}}$ are chosen, the performance of the models is similar (shown in Appendix A).
Figure 6 (b) shows the MSE for the first TDM $\|\bm{\mu}{0,S_{1}}\|_{2}$. Here, the overall discrepancy between quantum and classical models is smaller than for $\Delta E_{T_{1}}$. Nevertheless, for larger training dataset sizes, the siQNN-NN still performs about half an order of magnitude better than the classical models, but the performance of the siQNN is similar to that of the classical models. For the other TDMs (shown in Appendix A), the siQNN and siQNN-NN perform similarly or worse than the classical models for small training dataset sizes; however, as the training dataset size increases, the siQNN-NN significantly outperforms the other models. This indicates that the siQNN alone is not expressive enough to account for the $R$ dependence. This can be alleviated by parametrizing the interatomic distance with a classical NN.
@@ III.3.2 $\mathrm{H}_{2}$
Some of the target functions (Figure 4), such as $\Delta E_{T_{1}}$, are relatively trivial, which is also reflected in the MSE scores of the models (shown in Appendix A). For these trivial target functions, the MSE for some classical models is already in the order of $10^{-5}$ with a training dataset size of four, and it decreases with increasing training dataset size down to about $10^{-7}$. The quantum-enhanced models perform about an order of magnitude worse than the classical models for these functions because the increased feature space adds additional complexity to the learning task, which introduces unnecessary complexity and therefore overfitting. Further, as explained in Section II.3, we stop training the (Q)NN-based models when $\mathcal{L}_{\rm t}=10^{-6}$ is reached. Therefore, if the test MSE of these models is close to or below $10^{-6}$, further training with a smaller $\mathcal{L}_{\rm t}$ could reduce the test MSE even further for these models, potentially aligning them with models that do not use a $\mathcal{L}_{\rm t}$ such as the kernel methods, i.e., SVR or GPR. Consequently, for target functions and training dataset sizes with MSE values close to or below $\mathcal{L}_{\rm t}$, comparisons between the (Q)NNs and the kernel methods are limited significance. This effect becomes more relevant for increasing training dataset size.
@@
Overall, the classical models outperform the quantum-enhanced models for relatively trivial target functions. However, for more complex target functions, the quantum-enhanced models outperform the classical models, as the additional information from the ground state can be leveraged to achieve better prediction performances.
@@ III.3.3 $\mathrm{H}_{4}$
In this work, we investigated the effectiveness of enhancing the performance of a machine learning model trained to predict excited-state properties of a molecule in different geometric configurations. We achieved this by training a QNN directly on the molecular ground state to extract information from the wave function. The proposed model consists of a symmetry-respecting QNN combined with an NN, with the observable measured on the quantum state consisting only of Pauli-$Z$ operators. The performance of the siQNN-NN and siQNN were significantly better than the performance of the models learning only on classical data for small training dataset sizes across all molecules.
@@
From these observations, we deduce that the siQNN effectively extracts valuable information from the ground state and provides a good estimate of various target functions in this study. For very small training datasets, the additional complexity of the siQNN-NN leads to overfitting which decreases the performance relative to the siQNN. As the training dataset size increases, the siQNN, however, struggles to describe the training data points to the given accuracy, and the additional degrees of freedom in the NN become valuable to account for the $R$ dependence of the observable.
@@
We note several limitations of our study. Firstly, we expect that for each target function and molecule there exist a number of training data points above which the models learning purely on the classical data reach sufficient test dataset accuracy on average and therefore the additional effort of preparing the quantum information is not reasonable anymore. This is because for all data points, the ground state needs to be provided, introducing a relevant overhead relative to models learning only on the classical data. Further, to enable our model to predict excited-state properties, training values of those properties need to be derived using another technique. Therefore, our model is relevant for molecules where the properties of interest for different geometric configurations are calculable but with significant effort, while the ground state is less computationally demanding. Given that excited-state properties are generally more complex than ground-state properties, there are several molecules and properties that meet this condition.
@@
Secondly, the substructure of the siQNN has similarities to the structure of a quantum convolutional neural network, which is known to be simulable. This suggests the possibility that the siQNN is efficiently classically simulable. However, the envisioned use case for the procedure described in this work is a scenario where molecular ground state simulations are performed on a quantum computer. Adding a comparatively small QNN as suggested here is then a relatively small overhead to obtain excited state properties on top of ground state properties. Still, it could be an interesting further research path to construct a purely classical model inspired by our quantum-enhanced model that could extract valuable information and investigate whether it can do this even more efficiently from the ground, for example, using classical shadows [33]. Finally, this study was conducted using exact ground states. Investigating the effect of imperfect simulations of molecular grounds states as input is an interesting research question.
@@
In conclusion, we propose a procedure to deriving excited-state properties in an efficient way. We deem this approach valuable to analyzing molecular properties and understanding and predicting their behaviors. We view this study as a starting point for further investigations into extracting information from the ground state for excited-state properties.
@@ Acknowledgements.
Appendix A MSE Score Plots

### v2

@@ III.3 Benchmark
To draw conclusions about the performance of the models based on training dataset size independent of the specific choice of points, we sample multiple training datasets $\mathcal{D}_{l}^{L}\subset\mathcal{D}$. Each set $\mathcal{D}_{l}^{L}$ contains $L$ data points. For each $L$, we sample 50 independent datasets, i.e., $l=1,\dots,50$. As a test set, we choose the remaining data points for each $l$ and $L$, i.e., $\mathcal{D}\setminus\mathcal{D}_{l}^{L}$. In the following, all MSE scores shown are derived from the test dataset. We aim to simulate a realistic scenario where sampling is expensive, and the approximate locations of three different regions (attractive, equilibrium, and repulsive) along the dissociation curve are known. From each of these regions one data point is randomly chosen to account for the different behaviors in these regions. These three data points are then used as initial points for a Bayesian optimization, which determines the final data points in the training dataset $\mathcal{D}_{l}^{L}$. However, using a less informed sampling method, e.g., equally spaced training data points from random subsets of the given dataset yields similar results.
@@ III.3.1 LiH
Figure 6 (a) exemplarily shows a boxplot of the MSE for various models for the first transition energy $\Delta E_{T_{1}}$. The median MSE on the test set for the quantum models (siQNN and siQNN-NN) are similar to each other and nearly an order of magnitude lower than those of the classical models (SVR, GPR, NN), which themselves perform similarly across all training dataset sizes considered. Additionally, the variation of the MSE for the siQNN and siQNN-NN is significantly larger than that of the classical models. This is because, for some training dataset samples, the training points poorly represent the overall shape of the target function, causing all models to perform equally poorly. For other training data point combinations, the siQNN and siQNN-NN outperform the classical models by about an order of magnitude in MSE. In cases where other transition energies than $\Delta E_{T_{1}}$ are chosen, the performance of the models is similar (shown in Appendix D).
Figure 6 (b) shows the MSE for the first TDM $\|\bm{\mu}{0,S_{1}}\|_{2}$. Here, the overall discrepancy between quantum and classical models is smaller than for $\Delta E_{T_{1}}$. Nevertheless, for larger training dataset sizes, the siQNN-NN still performs about half an order of magnitude better than the classical models, but the performance of the siQNN is similar to that of the classical models. For the other TDMs (shown in Appendix D), the siQNN and siQNN-NN perform similarly or worse than the classical models for small training dataset sizes; however, as the training dataset size increases, the siQNN-NN significantly outperforms the other models. This indicates that the siQNN alone is not expressive enough to account for the $R$ dependency and potentially to find the exact eigenbasis. This can be alleviated by parameterizing the interatomic distance with a classical NN.
@@ III.3.2 $\mathrm{H}_{2}$
Some of the target functions (Figure 4), such as $\Delta E_{T_{1}}$, are relatively trivial, which is also reflected in the MSE scores of the models (shown in Appendix D). For these trivial target functions, the MSE for some classical models is already in the order of $10^{-5}$ with a training dataset size of four, and it decreases with increasing training dataset size down to about $10^{-7}$. The quantum-enhanced models perform about an order of magnitude worse than the classical models for these functions because the increased feature space adds additional complexity to the learning task, which introduces unnecessary complexity and therefore overfitting. Further, as explained in Section II.3, we stop the training of the (Q)NN-based models when $\mathcal{L}_{\rm t}=10^{-6}$ is reached. Therefore, if the test MSE of these models is close to or below $10^{-6}$, further training with a smaller $\mathcal{L}_{\rm t}$ could reduce the test MSE even further for these models, potentially aligning them with models that do not use a $\mathcal{L}_{\rm t}$ such as the kernel methods, i.e., SVR or GPR. Consequently, for target functions and training dataset sizes with MSE values close to or below $\mathcal{L}_{\rm t}$, comparisons between the (Q)NNs and the kernel methods are of limited significance. This effect becomes more relevant for increasing training dataset size.
@@ III.3.3 $\mathrm{H}_{4}$
Overall, one can observe that the classical models outperform the quantum-enhanced models for relatively trivial target functions already for the considered low data regime. However, for more complex target functions, the quantum-enhanced models outperform the classical models, as the additional information from the ground state can be leveraged to achieve better prediction performances.
@@
In this work, we investigated the effectiveness of enhancing the performance of a machine learning model trained to predict excited-state properties of a molecule in different geometric configurations. We achieved this by training a QNN directly on the molecular ground state to extract information from the wave function. The proposed model consists of a symmetry-respecting QNN combined with an NN, with the observable measured on the quantum state consisting only of commuting Pauli-strings. Pauli-$Z$ operators. The performance of the siQNN-NN and siQNN were significantly better than the performance of the models learning only on classical data for small training dataset sizes across all molecules with a test MSE reduction of up to two orders of magnitude.
@@
From these observations, we deduce that the siQNN effectively extracts valuable information from the ground state and provides a good estimate of various target functions in this study. For very small training datasets, the additional complexity of the siQNN-NN leads to overfitting which decreases the performance relative to the siQNN. As the training dataset size increases, the siQNN, however, struggles to describe the training data points to the given accuracy, and the additional degrees of freedom in the NN become valuable to account for the $R$ dependence and the potentially imperfect preparation of the eigenbasis of the target observable with the siQNN. With increasing data set size, the performance gap between classical and quantum models closes. We observe that the exact number of data points where this can be observed is dependent on the complexity of the target function, that is, more complex target functions require more data points for classical models to be competitive with quantum models.
@@
We note several limitations of our study. Firstly, we expect that for each target function and molecule there exist a number of training data points (depending on its complexity) above which the models learning purely on the classical data reach sufficient test dataset accuracy on average and therefore the additional effort of preparing the quantum information is not reasonable anymore. This is because for all data points, the ground state needs to be provided, introducing a relevant overhead relative to models learning only on the classical data. Further, to enable our model to predict excited-state properties, training values of those properties need to be derived using another technique. Therefore, our model is relevant for molecules where the properties of interest for different geometric configurations are calculable but with significant effort, while the ground state is less computationally demanding. Given that excited-state properties are generally more complex than ground-state properties, there are several molecules and properties that meet this condition.
@@
Secondly, the substructure of the siQNN has similarities to the structure of a quantum convolutional neural network, which is known to be simulatable. This suggests the possibility that the siQNN is efficiently classically simulatable. However, the envisioned use case for the procedure described in this work is a scenario where molecular ground state simulations are performed on a quantum computer. Adding a comparatively small QNN as suggested here is then a relatively small overhead to obtain excited state properties on top of ground state properties. Still, it could be an interesting further research path to construct a purely classical model inspired by our quantum-enhanced model that could extract valuable information and investigate whether it can do this even more efficiently from the ground, for example, using classical shadows [38].
Finally, this study was conducted using exact ground states. Investigating the effect of imperfect simulations of molecular ground states as input is an interesting research question. In Appendix B we further estimate the effect of shot noise on the performance of our machine learning approach and find promising results. Regarding decoherence and gate noise, given that using the circuit in the paper extends the circuit depth only logarithmically in the number of qubits compared to the preparation of the ground state, we have good reason to believe that our method does not significantly deviate from results obtained from NISQ state-preparation circuits (e.g. VQE). More specifically, the results will be comparable to other types of QNNs of similar size (see e.g. Ref. [39], which has used a similar observable on real quantum computers). A more rigorous investigation into the effect of shot noise and additionally the effect of other noise types apparent in current quantum computers could be conducted in follow-up work.
In order to estimate the scaling of our model to larger molecules, we confirmed its applicability and performance advantages with different system sizes, namely four, five, and six orbitals. However, these are small problem sizes compared to many molecules of interest, such as iron complexes, which require too many qubits to be prepared on today’s quantum simulators and devices. A potential problem for these larger system sizes could be the complexity of finding the eigenbasis of the target observable. A potential approach could be to prepare multiple copies of the ground state followed by QNNs, whose measurement is added up and, therefore, optimized in parallel, so that the different copies could prepare different common eigenstates of commuting subsets of the target observable, to break down the eigenbasis search task.
@@
In conclusion, we propose a procedure to derive excited-state properties in an efficient way. We consider this approach valuable to analyzing molecular properties and understanding and predicting their behaviors. We view this study as a starting point for further investigations into extracting information from the ground state for excited-state properties.
@@ Acknowledgements.
Appendix A QNN Parameter Scaling
In this section, we show that the parameters in the QNN scale linearly with the number of orbitals $n_{\rm orb}$, i.e, $n_{\mathrm{params}}=8(n_{\mathrm{orb}}-1)-3$. For this we need to note that the first layer consists of $n_{\mathrm{orb}}$ gates, which we call $N$ here, each of which contain three parameters. The subsequent pooling layer consists of $\frac{n_{\mathrm{orb}}}{2}$ gates, called $P$ here, with two parameters each. The next set of $N$ and $P$ gates act on half the qubits and therefore their number of parameters is half the number of parameters of the first layer and so on. In the last layer the circular application of the $N$ gate is not necessary since only two qubits are left each and therefore there are three parameters less in the last layer of $N$ and $P$ gates, which otherwise would consist of eight parameters.
$n_{\mathrm{params}}+3$ $=3n_{\mathrm{orb}}+n_{\mathrm{orb}}+3\frac{n_{\mathrm{orb}}}{2}+\frac{n_{\mathrm{orb}}}{2}+...+8$ (9)
$=4n_{\mathrm{orb}}+2n_{\mathrm{orb}}+...+8$ (10)
$=\sum_{i=1}^{\log_{2}(n_{\mathrm{orb}})}2^{i+2}=4((\sum_{i=0}^{\log_{2}(n_{\mathrm{orb}})}2^{i})-1)$ (11)
$=4(\frac{2^{\log_{2}(n_{\mathrm{orb}})+1}-1}{2-1}-1)=4(2n_{\mathrm{orb}}-2)$ (12)
$=8(n_{\mathrm{orb}}-1)$ (13)
We do not need to pool down to two qubits left in the circuit but can measure earlier and therefore this is in general an upper limit on the number of parameters in the circuit.
Figure 8: The influence of shot noise on the siQNN-NN for a set of training data points and the target function $\Delta E_{T_{2}}$ for LiH. In (a) the predicted test and train energy differences is shown after training and evaluation with 10,000 shots. In (b) the test MSE depending on the number of shots used for training and evaluation for the siQNN-NN is shown. The “evaluation 100k” shows the performance when evaluating instead with 100,000 shots. As comparison the MSE on the measurement of the Hamiltonian on the ground state is shown. Here ”upper bound” means that each Pauli-string is measured with the given number of shots while “qwc” means that the given shot budget is shared across qubit wise commuting subsets of the Pauli-strings in the Hamiltonian.
Appendix B Influence of Shot Noise
In current quantum hardware several kinds of noise need to be considered. Due to error correction in the upcoming years the influence of most noise types can be significantly reduced, but shot noise is a challenge intrinsic to quantum computers and will therefore persist beyond the NISQ regime. Therefore in this section we focus on this type of noise and investigate its effects. One of the advantages of the siQNN model is that the number of required shots for a given accuracy is expected to be significantly reduced compared to conventional methods such as VQE extensions. This is because the model directly approximates $\Delta E$ on the quantum state. In order to investigate into this shot reduction we train and test our model for the target function and training data points from Figure 5(a) for different numbers of shots. Figure 8(a) exemplary shows the prediction of the siQNN-NN after training and evaluation with 10.000 shots. We can see that in this case there is no systematic error introduced by the shot noise and that already with $\sim 10^{4}$ shots the test MSE is lower then the test MSE of the classical comparison models [compare Figure 5(a)].
In Figure 8(b) we show the resulting test MSE’s with the given number of shots for the training and test evaluations with the siQNN-NN. The label “siQNN-NN, evaluation 100k” shows the test MSE when training with the given number of shots and evaluating with 100,000 shots. We show this to differentiate the effect of noise on the training and the evaluation process. This scenario acknowledges that training requires a significantly larger number of model evaluation than inference such that a larger shot budget can be granted at inference.
We compare these results to measuring the exact Hamiltonian (H) on the ground state with different numbers of shots. From this only the ground state energy is derived. To calculate $\Delta E$, the excited state energy would be needed, too. This would only increase the error on $\Delta E$. We, however, neglect this contribution here such that the resulting error can be interpreted as a lower bound for determining $\Delta E$ directly through measurement. The ”upper bound” label shows the MSE when measuring each Pauli-string in the Hamiltonian with the given number of shots. In contrast to the observable measured in the siQNN-NN the Hamiltonian consists of many non-commuting Pauli-strings and therefore can not be measured on the same quantum state. In the case shown here the Hamiltonian consists of 276 different Pauli-strings, which can be grouped to 58 qubit wise commuting (qwc) subsets, that can be measured on the same quantum state each. Therefore the total number of shots needs to be shared between these 58 sets, which reduces the number of shots available for each set significantly. The resulting MSE is labeled “qwc”. More elaborate grouping strategies can be leveraged such as from the classical shadows mechanism to increase the precision but the ”upper bound” case is an by far unreachable upper bound to this. Furthermore, many of today’s quantum computers use qwc by default.
In this plot we can see that for the given target function and training data points the siQNN-NN derives $\Delta E$ with a significantly lower test MSE then if one would measure with the Hamiltonian on the ground and excited state and subtract to derive $\Delta E$. Even if we neglect the derivation of the excited state energy and the fact that the Hamiltonian consists of many non-commuting Pauli-strings, we need about 100,000 shots to derive $\Delta E$, whereas the siQNN-NN only needs about 30,000 shots in total to achieve a similar accuracy. In a more realistic scenario where the fact that we can only measure commuting sets on the same quantum state is considered we would need about 6 million (a factor of 58) shots for the same accuracy. Therefore our model reduces the number of shots necessary to derive $\Delta E$ for a given accuracy by more then two orders of magnitude relative to traditional methods for the given target function and training data set. Further we can see (“evaluation 100k”) that due to the shot noise in the training process a varying offset is introduced because of the few training data points considered so that the advantage of sampling the test data with a larger number of shots is not significant.
Figure 9: Average rank of the models, broken down by training dataset size for LiH, with all models equal between (a), (b), and (c) except the NN. The number of parameters in the NN corresponds in (a) to 70, in (b) to 150, and in (c) to 1000. Each point represents the relative rank of the respective model compared to all other models averaged over all considered targets and training samples for a molecule. The error bars indicate the standard errors of the mean values of those mean rankings.
Appendix C Larger NN comparison
The small size of the considered NN serves two purposes. Firstly, we want to derive the effect of the NN in the siQNN-NN and therefore need a model with similar size. Secondly, in this low data regime, we consider, one needs to be especially aware of overfitting, which can be reduced by restricting the expressivity of the model (in the example of the NN by reducing its size). In Figure 9 we can see that with an increasing number of parameters in the NN the performance of the NN worsens due to overfitting. We can see that for extremely few training data points (four to five) the effect of overfitting is less relevant. This is due to the fact that in this case many functions can fit the training points with low effort, and no overfitting appears in the pursuit to reach arbitrary precision on the training set.
Appendix D MSE Score Plots